OpenStudy (anonymous):

lim x→−6− tan (πx/4) =

OpenStudy (agreene):

\[\infty\]

OpenStudy (zarkon):

\[\lim_{x\to-6^{-}}\tan(\pi x/4)=\infty\]

OpenStudy (anonymous):

Thanks a ton!